The density of mercury is 13.6. What is the value of b?

Grade:12

1 Answers

Vikas TU
7 years ago
This has already been answered by one of our member:
Reposting the answer
The total number of atoms in 1 g of Hg are
 
 
= 6.023 *1023/200 = 6.023 *1023/2 *102
= 3.0115 *1021 = 3.012 * 1021
Density of Mercury (Hg) = 13.6 g/c.c.
 mass of 3.012 * 1021 atoms = 1/3.012 *1021
Now volume of 1 atom of mercury (Hg)
= 1/3.012 *1021 *13.6 c.c. = 103 *10/3012 *1021 * 136 c.c.
= 10-17/3012 *136 c.c. 10-17/409632 c.c. = 1000000 *10-23/409632 c.c.
 
 
= 2.44 *1023 c.c.
Since each mercury atom occupies a cube of edge length equal to its diameter, therefore,
Diameter of one Hg atom = (2.44 * 10-23)1/3 cm
= (24.4 * 10-24)1/3 cm.
= 2.905 * 10-8 cm ≡ 2.91